OpenStudy (anonymous):

please help, i am really confuse write the equation of the line thorough (2,-3) and m=0

OpenStudy (anonymous):

I assume you want slope-intercept form (y=mx+b), where m is equal to the slope, (x,y) is a coordinate on the line, and b is a random constant. "Plug-in" the numbers to get -3=0(2)+b. Simplify to get -3=b. Now take your m and b and put it into the original equation (y=mx+b) to get y=0x-3 or y=-3

OpenStudy (anonymous):

thanks :)